# Affective Computing in the Era of Large Language Models: A Survey from the NLP Perspective

### v1

@@ 4.1 Affective Understanding
In addressing affective computing tasks, PLMs typically necessitate separate approaches for affective understanding and generation tasks. In contrast, LLMs can unify all affective computing tasks into a sequence generation format. However, directly employing LLMs for AU tasks does not offer significant advantages over fine-tuned PLMs. Consequently, numerous recent studies focuse on constructing task-specific instruction tuning templates tailored to the unique characteristics of various tasks. These studies apply instruction tuning to the models, aiming to enhance the performance of LLMs on specific tasks or within particular domains.
4.1.1 Polarity & Emotional Classification
@@
Polarity Classification (PC) and Emotional Classification (EC) tasks are frequently employed to assess a model’s comprehension and perceptual abilities. Several studies pioneer the exploration of the effects of instruction tuning of LLMs on the SA task [122, 123]. The T5[124] series model, in particular, demonstrates an improved understanding of the sentiment content of text after instruction tuning. Recent research [123] focuses on exploring the different PEFT methods, adopting the LoRA and P-tuning methods to fine-tuning ChatGLM2[125]. Experimental results demonstrate that after instruction tuning, LLMs are greatly improved in eight datasets compared with the model without instruction tuning. However, the experimental results must conclusively determine which instruction tuning method is more effective in enhancing performance on PC. The emotional polarity of financial data is also a PC task. A recent study proposes a retrieval-augmented LLMs framework [126] consisting of instruction-tuned LLMs and a retrieval-augmented component. This method can classify the sentiment of data in the financial domain.
While these studies demonstrate LLMs’ effectiveness in PC and EC tasks, it is equally important to understand their underlying sentiment processing mechanisms. Research shown that words with sentiment information can influence the overall sentiment of the text [38]. Building on this understanding, LLMs leverage techniques of instruction tuning to grasp word polarity, enabling them to effectively perform methods PC and EC tasks.
Although these studies primarily focus on text-based PC and EC tasks, it is essential to recognize that real-world user expressions are not limited to text alone. User-generated content often encompasses multiple modalities, including images and audio, which can provide additional context and nuance to sentiment analysis. WisdoM [39] leverages contextual world knowledge derived from LMM to enhance multimodal sentiment analysis (MSA). Additionally, the authors introduce a training-free module called "contextual fusion," which can reduce noise in the contextual data and improve model classification accuracy. WisdoM significantly outperforms existing state-of-the-art methods on the MSED dataset, demonstrating its effectiveness in integrating contextual knowledge to improve emotion classification. Despite these promising advancements, it is worth noting that research on MSA tasks based on LLMs or Multimodal Large Language Models (MLLMs) still needs to be improved.
@@ 4.1.2 Aspect-Based Sentiment Analysis
Aspect-Based Sentiment Analysis (ABSA) encompasses multiple subtasks, such as Aspect Term Extraction (ATE), Aspect-oriented Sentiment Classification (ASC), Joint Aspect-Sentiment Analysis (JASA), and Aspect-Category-Opinion-Sentiment (ACOS), and others. The difficulty of these subtasks progressively increases, starting with known aspect items, then identifying aspect items, and finally performing the corresponding sentiment analysis. Traditional PLMs often design a separate module for each ABSA subtask, which reduces the correlation between ABSA subtasks. LLMs can leverage their strong generalization abilities to construct a unified framework for all ABSA subtasks.
[…]
4.1.3 Emotion Recogniztion in Conversation
Emotion Recogniztion in Conversation (ERC) tasks involve recognizing the sentiment label of each utterance in the conversation, which is highly dependent on the conversation history and the speaker.
@@ 4.1.3 Emotion Recogniztion in Conversation
InstructERC addresses these challenges by developing a template that includes a system prompt, dialogue history and the label map. It also introduces Instruction Tuning using LoRA for ERC tasks [49]. Additionally, InstructERC creates two auxiliary tasks, speaker identification and emotion impact prediction, to enhance the model’s understanding of speaker information. Building on this approach, CKERC [48] also employs LoRA to solve ERC tasks. It further leverages the LLaMA2-chat model [130] to generate common sense knowledge, enabling a more comprehensive understanding of the model of dialogue context. This model achieves state-of-the-art results on the MELD [131] and EmoryNLP [132] datasets.
@@
Researchers extend ERC tasks to incorporate multimodal inputs, considering the frequently multimodal aspects of human communication, leading to the development of Multimodal Emotion Recognition in Conversation (MERC). DialogueLLM [133] generates text descriptions of visual information through GPT-4, combines these descriptions with the conversation history as an instruction dataset, and uses LoRA to train LLaMA2. Through experiments, it has been found that adding information from other modalities can improve the model’s ability to understand dialogue information.
Due to the high similarity of mission objectives and approach between the SA and ERC tasks, researchers try to build a unified model that can accomplish both tasks simultaneously. Emollms [134] develops a multi-task instruction dataset, AAID, covering SA and ERC. Experimental results show improved performance on each task after multi-task training. Similarly, Unimse [47] introduces a multimodal affective knowledge-sharing framework, unifying MSA and MERC tasks and fusing information from multiple modalities through the T5 model. Notably, it is the first method to solve MSA tasks in a generative method. These studies indicate that mixing datasets from multiple tasks can improve a model’s generalization ability, enabling performance on untrained datasets comparable to fine-tuned PLMs.
@@ 4.1.4 Emotion Cause Extraction
Emotion Cause Extraction (ECE) has garnered widespread attention in recent years. This task aims to extract potential causes behind certain emotions in text, but its application scenarios are significantly limited due to the need for annotating emotional information. To address this issue, the Emotion-Cause Pair Extraction (ECPE) task propose. The ECPE task simultaneously extracts potential Emotion-Cause pairs and their corresponding causes within a document, thereby reducing the dependence on emotion labels [135].
Most current research on ECPE tasks involves fine-tuning traditional PLMs [136, 137] or leveraging the understanding capabilities of LLMs using well-designed prompts [50]. These methods primarily focus on the Textual Emotion-Cause Pair Extraction in Conversations task (TECPE). However, emotion causes may also be embedded in other modalities, leading to the development of Multimodal Emotion-Cause Pair Extraction in Conversations task (MECPE) [138]. LLMs shows potential in better integrating information from multiple modalities for such tasks.
To address the MECPE task, NUS-Emo [139] proposes a dual-component system that operates in two stages. In the first stage, the system uses ERC data to fine-tune the LLMs, enhancing its emotion recognition capability. It then employs ImageBind to encode image information fused with the recognized emotional information. In the second stage, the system inputs the emotional information into the model for inference and uses the inferred dataset as the training set for further fine-tuning. This iterative fine-tuning approach aims to improve the model’s ability to infer emotion causes across multiple modalities.
@@ 4.1.5 Subjective Text Analysis
LLMs are employed to explore the Stance Classification task. In this task, the model determines whether a user agrees, disagrees, or is neutral on a given point of view [140]. The study investigates ten open-source models and seven prompting schemes, revealing exciting conclusions. These include the fact that LLMs are competitive with in-domain supervised models and that the fine-tuning process only sometimes leads to better performance than zero-shot models. This is likely because the fine-tuning makes the model too specialized, thus unable to generalize to out-of-domain data points and varying stance definitions.
@@ 4.2 Affective Generation
In the current landscape of artificial intelligence, content generation that expresses sentiment is applied to various practical fields, including emotional dialogue, chatbots, emotional support conversation, and more [5]. It enhances humanized services by analyzing and generating emotional information from users.
@@ 4.2.1 Emotional Dialogue
Emotional Dialogue (ED) refers to the ability to understand and share another person’s feelings and then respond in a way that conveys this understanding and compassion. LLMs must to possess high emotional intelligence to resonate with users in their replies. However, due to the datasets’ limited and relatively homogeneous nature, the generated responses often tend to be formulaic and mechanical, resulting in lower-quality interactions. Researchers aim to construct more diverse empathetic dialogue datasets. Considering the high labor costs of manually constructing datasets, current research primarily employs synthesis data for LLMs methods to build high-quality empathetic dialogue datasets [141, 53, 55].
Several studies use these methods to create diverse datasets. For example, the ExTES dataset is expanded using GPT-3.5-turbo and self-instruct [142], resulting in a total of 11,177 conversations where users display negative emotions and the assistant provides more positive support responses [141]. Soulchat [58] employs GPT-3.5-turbo to expand existing dialogue datasets through iterative rounds, using prompts to generate multi-turn conversations. Another notable example is STICKERCONV, a multi-scene, multimodal empathetic dialogue dataset built by generating user information from a multi-agent system [55]. Data generation through LLMs can significantly reduce labor costs. The dialogue scenarios and content are wider than providing more diverse data types. However, the quality of the generated data depends on the model’s capabilities and may need to be more consistent, often requiring secondary screening. To mitigate quality issues, some studies avoid the direct generation of dialogue data and instead generate objective knowledge, thus avoiding the quality problem of the generated data. A recent study [53] proposes a novel approach using GPT-3.5-turbo to generate common sense knowledge about conversations. This generated knowledge encompasses the discourse’s underlying reasons, emotions, and intentions. By incorporating this knowledge into the model’s training process, the research enhances the model’s ability to comprehend the deeper meaning of conversations and generate more empathetic responses. Experimental results demonstrate that each type of knowledge is adequate for empathetic response generation.
@@ 4.2.2 Chatbot
Current datasets still limit most of the ERG and ESC tasks described above and need to capture the diversity of users and scenarios in real life fully. Chatbots capable of understanding user input and emotions are designed to address the practical application of emotional dialogue systems. These chatbots can offer encouragement during adversity and share in users’ happiness.
PICA [60] and EMOLLM [59] are two notable examples of such chatbots. While LLMs can function well as "assistants," their emotional intelligence is often lacking, and long replies frequently fail to convey empathy. To address this issue, PICA is proposed. It uses P-tuning to fine-tune the ChatGLM2-6B model on two synthetic Chinese empathetic dialogue datasets, effectively enhancing the emotional capabilities of LLMs. EMOLLM integrates a series of LLMs to support and assist users with mental health counseling tasks. It uses the QLoRA[143] method to fine-tune multiple LLMs, such as Qwen-7B, LLaMA3-8B-Instruct, and others, across different languages. By supporting various roles, EMOLLM enables the chatbot to generate more contextually appropriate user responses.
@@ 4.3 Multi-task of Affective Computing
However, current fine-tuning methods present limitations when dataset sources differ across tasks. Fine-tuning a model under these conditions can create conflicts and impact the original world knowledge embedded in the model [144]. These challenges highlight the need for more research on ensuring high-quality performance across all sentiment analysis tasks.
To address these issues, researchers propose a novel Modular Emotional Intelligence enhancement method (MoEI) featuring two collaborative techniques—Modular Parameter Expansion (MPE) and Intra-Inter Modulation (I2M)—is proposed to fit most emotion analysis tasks [145]. In conjunction with this method, the study develops EIBENCH, a unified affective computing dataset containing 88 datasets across 15 EI-related tasks, is introduced. MoEI adopts the Mixture-of-LoRA (MoLoRA) method for fine-tuning to improve EI while maintaining the generative ability of the model.
In multimodality, MEILLM[146] integrates deep feature vectors extracted from visual, textual, and audio models into baichuan13B-chat through the MLP layer. After fine-tuning the instructions, he can complete a variety of emotional interaction tasks, including psychological assessment, psychological portrait, psychological counseling, etc.

### v2

@@ 4.1 Affective Understanding
In addressing affective computing tasks, PLMs typically necessitate separate approaches for affective understanding and generation tasks. In contrast, LLMs can unify all affective computing tasks into a sequence generation format. However, directly employing LLMs for AU tasks does not offer significant advantages over fine-tuned PLMs. Consequently, numerous recent studies focus on constructing task-specific instruction tuning templates tailored to the unique characteristics of various tasks. These studies apply instruction tuning to the models, aiming to enhance the performance of LLMs on specific tasks or within particular domains.
4.1.1 Polarity and Emotional Classification
@@
Polarity Classification (PC) and Emotional Classification (EC) tasks are frequently employed to assess a model’s comprehension and perceptual abilities. Several studies pioneer the exploration of the effects of instruction tuning of LLMs on the SA task Raffel et al. (2020b); Peng et al. (2024). The T5 Raffel et al. (2020a) series model, in particular, demonstrates an improved understanding of the sentiment content of text after instruction tuning. Recent research Peng et al. (2024) focuses on exploring the different PEFT methods, adopting the LoRA and P-tuning methods to fine-tuning ChatGLM2 GLM et al. (2024). Experimental results demonstrate that after instruction tuning, LLMs are greatly improved in eight datasets compared with the model without instruction tuning. However, the experimental results must conclusively determine which instruction tuning method is more effective in enhancing performance on PC.
While these studies demonstrate LLMs’ effectiveness in PC and EC tasks, it is equally important to understand their underlying sentiment processing mechanisms. Research shown that words with sentiment information can influence the overall sentiment of the text Gan et al. (2023). Building on this understanding, LLMs leverage techniques of instruction tuning to grasp word polarity, enabling them to effectively perform methods PC and EC tasks. These studies primarily focus on text-based PC and EC tasks, it is essential to recognize that real-world user expressions are not limited to text alone. User-generated content often encompasses multiple modalities, including images and audio, which can provide additional context and nuance to sentiment analysis. WisdoM Wang et al. (2024f) leverages contextual world knowledge derived from MLLM to enhance multimodal sentiment analysis (MSA). Additionally, the authors introduce a training-free module called "contextual fusion", which can reduce noise in the contextual data and improve model classification accuracy. WisdoM significantly outperforms existing state-of-the-art methods on the MSED dataset, demonstrating its effectiveness in integrating contextual knowledge to improve emotion classification. In WisdoM, the method of converting images into text captions results in the loss of a significant amount of image information and fails to focus on the more detailed emotional expressions within the images. EMO-LLaMA Xing et al. (2024) addresses this issue by training a dedicated encoder for facial features to extract both global and local information from faces. It employs handcrafted prompts to characterize user profiles and combines the previously extracted facial information to fine-tune the LLaMA model, thereby alleviating the model’s inaccuracies in recognizing image details. This model achieved the highest accuracy on both static and dynamic facial expression recognition datasets. Emotion-LLaMA Cheng et al. (2024b) additionally encodes audio; it uses a specific emotional encoder to seamlessly integrate audio, video, and text modalities, and performs LoRA fine-tuning on the LLaMA2 model. It achieved the highest score on the EMER dataset. Omni-Emotion Yang et al. (2025) has a similar model structure to Emotion-LLaMA and explores various fusion schemes of facial features with other visual features. Experiments found that the method of separately extracting the visual and facial features of the entire video and then concatenating them has the least impact on MLLMs.
@@ 4.1.2 Aspect-Based Sentiment Analysis
Aspect-Based Sentiment Analysis (ABSA) encompasses multiple subtasks, such as Aspect Term Extraction (ATE), Aspect-oriented Sentiment Classification (ASC), Joint Aspect-Sentiment Analysis (JASA), and Aspect-Category-Opinion-Sentiment (ACOS), and others. Traditional PLMs often design a separate module for each ABSA subtask, which reduces the correlation between ABSA subtasks. LLMs can leverage their strong generalization abilities to construct a unified framework for all ABSA subtasks.
Recent research explore this potential. InstructABSA introduces Instruction tuning based on the Tkk-Instruct model Wang et al. (2022) for ABSA. This study finds that LLMs are sensitive to different instructions, with experimental results fluctuating by up to 10% Scaria et al. (2023). ITSCL Zhang et al. (2024b) is based on the T5 model and performs unified encoding of emotions and opinions. It introduces contrastive learning loss at each layer of the model to differentiate the feature vectors of different emotions and entity labels, thereby enhancing the model’s performance. In addition to strengthening the correlations between ABSA tasks, large language models can leverage their extensive world knowledge to enhance information exchange across different domains. Some reseraches Ding et al. (2024) propose an ABSA cross-domain learning fine-tuning framework. In the first stage, domain knowledge is used to train the domain variation adapter, while replay data is used to train the domain-invariant adapter, employing orthogonal constraints to force the model to learn the differences between domain-invariant and domain-variant knowledge. In the second stage, all replay data will be used to train the domain-invariant adapter while freezing the corresponding domain variation adapter. This two-stage fine-tuning effectively utilizes knowledge from different domains and avoids the issue of forgetting previous domain knowledge. The model achieved the best performance across 19 datasets.
The ABSA task shares similar objectives with sentiment analysis and sentiment polarity analysis. Unisa aims to build a unified multimodal emotion analysis model that integrates various sentiment analysis tasks Li et al. (2023b). This model has already established a new dataset and benchmark, SAEval, to facilitate research in this field. SAEval encompasses various ABSA sub-tasks, providing a comprehensive evaluation platform for future research.
4.1.3 Emotion Recognition in Conversation
@@
Emotion Recognition in Conversation (ERC) tasks involve recognizing the sentiment label of each utterance in the conversation, which is highly dependent on the conversation history and the speaker. InstructERC addresses these challenges by developing a template that includes a system prompt, dialogue history, and the label map. It also introduces Instruction Tuning using LoRA for ERC tasks Lei et al. (2024). Additionally, InstructERC creates two auxiliary tasks, speaker identification and emotion impact prediction, to enhance the model’s understanding of speaker information. This model performs very well on the MELD Poria et al. (2019a) and EmoryNLP Zahiri & Choi (2018) datasets.
@@
Researchers extend ERC tasks to incorporate multimodal inputs, considering the frequently multimodal aspects of human communication, leading to the development of Multimodal Emotion Recognition in Conversation (MERC). DialogueLLM Zhang et al. (2023h) generates text descriptions of visual information through GPT-4, combines these descriptions with the conversation history as an instruction dataset, and uses LoRA to train LLaMA2. Through experiments, it has been found that adding information from other modalities can improve the model’s ability to understand dialogue information.
@@ 4.1.4 Emotion Cause Extraction
Emotion Cause Extraction (ECE) has garnered widespread attention in recent years. This task aims to extract potential causes behind certain emotions in text, but its application scenarios are significantly limited due to the need for annotating emotional information. To address this issue, the Emotion-Cause Pair Extraction (ECPE) task propose. The ECPE task simultaneously extracts potential Emotion-Cause pairs and their corresponding causes within a document, thereby reducing the dependence on emotion labels Xia & Ding (2019b). Most current research on ECPE tasks involves fine-tuning traditional PLMs Ding et al. (2020); Wei et al. (2020b) or leveraging the understanding capabilities of LLMs using well-designed prompts Wu et al. (2024b). These methods primarily focus on the Textual Emotion-Cause Pair Extraction in Conversations task (TECPE). However, emotion causes may also be embedded in other modalities, leading to the development of Multimodal Emotion-Cause Pair Extraction in Conversations task (MECPE) Wang et al. (2024c). LLMs shows potential in better integrating information from multiple modalities for such tasks.
Recent work on multimodal emotion-cause understanding has progressed from clause extraction to explicit cause generation. NUS-Emo Luo et al. (2024) attacks the original MECPE task with a two-stage self-training pipeline: LLMs are first emotion-tuned on ERC data, ImageBind aligns visual and textual affective representations, and the resulting multimodal predictions are recycled as pseudo-labels for further fine-tuning, yielding steady gains in cross-modal cause inference. Moving beyond extraction, Wang et al. introduce the Multimodal Emotion Cause Generation in Conversations (MECGC) task and the ObG framework Wang et al. (2024b); an emotion-cause-aware captioner (ECCap) is trained to verbalise cause-focused video descriptions, which are then fed to an encoder–decoder generator, demonstrating that caption-based bridging markedly improves reason generation. FAME-Net Wang et al. (2024e) refines this idea by fusing holistic scene features from a visual encoder with fine-grained facial cues from VGG16, allowing dual-granularity guidance that localises both contextual and expressive triggers. Finally, Emotion-LLaMA Cheng et al. (2024b) instruction-tunes LLaMA with a dedicated “[Reason]” token so that emotion prediction and textual justification are produced jointly, enhancing model confidence and interpretability. Collectively, these studies highlight the effectiveness of ERC pre-training, caption-based intermediaries, multi-stream visual fusion, and explanation-oriented instruction tuning for multimodal emotion-cause modelling.
@@ 4.1.5 Subjective Text Analysis
LLMs are employed to explore the Stance Classification task. In this task, the model determines whether a user agrees, disagrees, or is neutral on a given point of view Cruickshank & Ng (2024). The study investigates ten open-source models and seven prompting schemes, revealing exciting conclusions. These include the fact that LLMs are competitive with in-domain supervised models and that the fine-tuning process only sometimes leads to better performance than zero-shot models. This is likely because the fine-tuning makes the model too specialized, thus unable to generalize to out-of-domain data points and varying stance definitions. In the opinion expression task, STOEI Jia et al. (2024) concatenates text embeddings and speech embeddings, inputting them together into the LLM for instruction fine-tuning, enabling the model to learn to extract opinion words from the input and analyze the sentiments associated with those opinion words. This approach surpasses existing techniques by over 9.20% and achieves state-of-the-art (SOTA) results.
Although LLMs possess inherent advantages in these subjective sentiment analysis tasks, the datasets for these tasks are mostly scarce, making instruction fine-tuning of the model challenging.
@@ 4.2 Affective Generation
In the current landscape of artificial intelligence, content generation that expresses sentiment is applied to various practical fields, including emotional dialogue, chatbots, emotional support conversation, and more Nie & Zhan (2022). It enhances humanized services by analyzing and generating emotional information from users.
@@ 4.2.1 Emotional Dialogue
Emotional Dialogue (ED) refers to the ability to understand and share another person’s feelings and then respond in a way that conveys this understanding and compassion. LLMs must to possess high emotional intelligence to resonate with users in their replies. However, due to the limitations of the dataset and its relative homogeneity, the generated responses often tend to be formulaic and mechanical, resulting in lower interaction quality. Some researchers aim to construct a more diverse empathetic dialogue dataset. Considering the high labor costs associated with manually constructing datasets, some current studies utilize LLM-generated synthetic data to build high-quality empathetic dialogue datasets.Zheng et al. (2023b); Wang et al. (2023b); Zhang et al. (2024f).
To enlarge training resources for empathetic dialogue, several works exploit LLM-driven self-augmentation. SoulChat Chen et al. (2023b) iteratively prompts GPT-3.5-turbo to grow multi-turn exchanges, yielding richer emotional trajectories. EKTC Cao et al. (2025) constructs the ED-TooL dataset by asking LLaMA3 to label the exact moments at which external tools should be invoked; the resulting high-quality tool-call exemplars are then used for instruction-fine-tuning, enabling models to decide autonomously—and more cleanly—when tool usage will enhance their responses. In the task of empathetic responses, the information dimension carried by a single text modality is relatively limited. Stickers and voice can convey non-verbal cues in interpersonal communication, and multimodal interaction mechanisms hold significant value in enhancing empathetic expression. StickerConv generates a multi-scenario, multimodal empathetic dialogue dataset constructed from user information through a multi-agent system, enabling more natural and fluid interactions with users by generating stickers within the dialogue Zhang et al. (2024f).
Speech–centric approaches further exploit prosody to boost empathy. BLSP-Emo Wang et al. (2024a) follows a two-step alignment strategy: (i) convert the speaker’s audio into an explicit textual emotion hint and let an LLM draft a reply; (ii) apply QLoRA to inject the raw audio encodings themselves, forcing the model to ground its responses directly in vocal affect. PerceptiveAgent Yan et al. (2024) transcribes prosodic features with an audio-encoder–text-decoder pair, feeds the resulting description together with dialogue history to an LLM, and finally vocalises the generated reply via a synthesiser—thus closing the speech-to-speech loop. EMOVA Chen et al. (2024a) generalises this pipeline to multimodal inputs and, crucially, lets the model emit style tokens that steer the synthesiser, yielding richer and more engaging emotional speech.
@@ 4.2.2 Chatbot
Current datasets still limit most of the ERG and ESC tasks described above and need to capture the diversity of users and scenarios in real life fully. Chatbots capable of understanding user input and emotions are designed to address the practical application of emotional dialogue systems. These chatbots can offer encouragement during adversity and share in users’ happiness. PICA Zhang et al. (2023i) and EMOLLM EmoLLM (2024) are two notable examples of such chatbots. While LLMs can function well as "assistants," their emotional intelligence is often lacking, and long replies frequently fail to convey empathy. To address this issue, PICA is proposed. It uses P-tuning to fine-tune the ChatGLM2-6B model on two synthetic Chinese empathetic dialogue datasets, effectively enhancing the emotional capabilities of LLMs. EmoLLM is a series of large mental health models that can support the mental health counseling link of understanding users, supporting users, and helping users. It uses the QLoRADettmers et al. (2024) method to fine-tune multiple LLMs, such as Deepseek-R1_14b_int4, LLaMA3-8B-Instruct, and others, across different languages. By supporting various roles, EMOLLM enables the chatbot to generate more contextually appropriate user responses.
@@ 4.3 Multi-task of Affective Computing
Due to the high similarity of mission objectives and approach in affective understanding tasks, researchers try to build a unified model that can accomplish tasks simultaneously. Emollms Liu et al. (2024c) develops a multi-task instruction dataset, AAID, covering SA and ERC. Experimental results show improved performance on each task after multi-task training. GSA-7B Hou et al. (2024) undergoes progressive LoRA training on three tasks, SA, ERC, and SD that are both related and increasingly challenging. It has been found that this hierarchical training approach can effectively improve the model’s performance across multiple tasks. In addition to classification tasks, M2SE Li et al. (2024a) also introduces tasks such as sentiment cause extraction, achieving a unified fine-tuning framework for multimodal and multitask scenarios. By sharing model parameters and facilitating knowledge transfer between tasks, it enhances the model’s ability to generalize in complex emotional contexts. This model is applicable to various emotion computation tasks.
However, current fine-tuning methods present limitations when dataset sources differ across tasks. Fine-tuning a model under these conditions can create conflicts and impact the original world knowledge embedded in the model Li et al. (2024b). These challenges highlight the need for more research on ensuring high-quality performance across all sentiment analysis tasks. To address these issues, researchers propose a novel Modular Emotional Intelligence enhancement method (MoEI) featuring two collaborative techniques—Modular Parameter Expansion (MPE) and Intra-Inter Modulation (I2M)—is proposed to fit most emotion analysis tasks Zhao et al. (2024b). In conjunction with this method, the study develops EIBENCH, a unified affective computing dataset containing 88 datasets across 15 EI-related tasks, is introduced. MoEI adopts the Mixture-of-LoRA (MoLoRA) method for fine-tuning to improve EI while maintaining the generative ability of the model.
In multimodality, MEILLMDong et al. (2024) integrates deep feature vectors extracted from visual, textual, and audio models into baichuan13B-chat through the MLP layer. After fine-tuning the instructions, he can complete a variety of emotional interaction tasks, including psychological assessment, psychological portrait, psychological counseling, etc. MODA Zhang et al. (2025c) develop a multimodal artificial intelligence that is both emotionally and intellectually capable. It addresses the issue of attention imbalance among different modalities in large models through a duplex attention alignment mechanism and a modular attention mask, allowing the model to focus more on fine-grained information from auxiliary modalities. In terms of cognitive and emotional understanding, MODA can accurately identify user intentions and emotional inclinations, demonstrating immense potential in the field of human-computer dialogue.